# A Mutual Information Perspective on Multiple Latent Variable Generative Models for Positive View Generation

### v1

@@ 4.3 Continuous Sampling
All our encoders are trained using Continuous Sampling, except for SimCLR, which follows the previous setup. Additionally, to compare overall training speed to standard data loading, we trained a ResNet-$18$ model with SimCLR for $20$ epochs on ImageNet-$100$, on $4$ NVIDIA A100-SXM4-40GB GPUs and different batch sizes ($32\times 4$, $64\times 4$, $128\times 4$, $256\times 4$). The experiment compares the standard PyTorch Paszke et al. (2019) loader, the efficient FFCV Leclerc et al. (2023) loader (both with $8$ workers), and the BigBiGan generator. Figure 5 displays our findings, reporting the mean GPU seconds per epoch, the $\text{CO}_{2}$ emissions rate, and the total $\text{CO}_{2}$ emissions estimated using CodeCarbon Schmidt et al. (2021). Continuous Sampling proves significantly faster than torch loader and only marginally slower than FFCV. In terms of $\text{CO}_{2}$ emissions rate, the use of BigBiGan led to higher energy consumption, due to intensive GPU usage. Nevertheless, in terms of total $\text{CO}_{2}$ the values remain comparable with torch. In conclusion, the increasing efficiency and precision of modern image generation models, especially fast-sampling GANs, make Continuous Sampling an interesting alternative to conventional data-loading techniques, allowing great image variability while maintaining comparable training times.
@@ 5 Discussion and Conclusions
Our work showcases MLVGMs as a distinct category of models, offering new tools to assess the impact of latent variables. Specifically, the proposed Monte Carlo quantification method supports previous empirical observations on the “global-to-local” nature of MLVGMs, but allows a more in-depth and quantitative analysis. As a result, we reveal that modern gan-based MLVGMs, such as BigGan and StyleGan employ over or under-utilized variables in the generative process, setting up a base for possible architectural improvements. In terms of view generation, our method has proven its superiority, surpassing previous perturbation strategies applied to single-variable models. However, it does not address the inherent challenge of SSCRL: views are defined upon “reasonable” thresholds, since “optimal” positives depend on the specific downstream task. Regarding generative models as a data source, they offer potential solutions to issues associated with real datasets, such as privacy concerns and usage rights Kaissis et al. (2020); DuMont Schütte et al. (2021). However, since generative models can inherit biases from the original data Asim et al. (2020), techniques to mitigate these biases could be considered Tan et al. (2020); Teo et al. (2023).
@@ Appendix B Training Dynamics
As explained in Section 3.1, we train each perturbation function $T_{\mathbf{z}}^{i}$ separately, and initialize as the identity function. This allows InfoNCE loss in Equation 5 to start from a very low value. During training, $T_{\mathbf{z}}^{i}$ progressively enhances diversity in the views, reducing $I(\mathbf{X};\mathbf{X}^{\prime})$, and causing InfoNCE to increase over time. An example of these training dynamics is illustrated in Figure 6(a). Figure 6(b) shows some qualitative examples of how views in the pixel space evolve during training. Specifically, the image on the left represents the anchor, while other images show the corresponding perturbed view as the magnitude increases.
@@ Appendix C Continuous Sampling
We detail the batch generation procedure for anchors and positive views in Algorithm 1. As explained in the main paper, the algorithm requires loading the pre-trained generative model (MLVGM in our case) on the same GPU device as the encoder model to be trained. In this way, the generated batch will be already on the correct device, eliminating the need for data loading.
@@
1:$\mathcal{A}\leftarrow\emptyset$triangleright initialize empty anchors set
2:$\mathcal{P}\leftarrow\emptyset$triangleright initialize empty positives set
@@
4:     $A_{b}\leftarrow\emptyset$triangleright initialize empty set of anchor variables
5:     $P_{b}\leftarrow\emptyset$triangleright initialize empty set of positive variables
@@
7:         $a_{i}\sim N^{t}(\mu,\sigma,t)$triangleright sample anchor code
8:         $p_{i}\leftarrow T_{\mathbf{z}}^{i}(a_{i})$triangleright generate positive code
@@
12:     $\mathbf{a}_{b}\leftarrow g(A_{b})$triangleright generate anchor image
13:     $\mathbf{p}_{b}\leftarrow g(P_{b})$triangleright generate positive image
@@
17:return $\mathcal{A},\mathcal{P}$triangleright SSCRL Encoder’s inputs
@@ Appendix D View Generation Training Details
For sampling positives, the random baselines are trained with $\mathcal{N}^{t}(0.0,0.2,2.0)$ and $\mathcal{N}^{t}(0.0,0.25,0.9)$ for each generator, respectively BigBiGan and StyleGan2. Following the suggestions of the original work (Li et al., 2022), we monitor the generated views and select the best walkers checkpoints for the learned baselines. For our ML views, BigBiGan samples positives by fixing the first latent and with parameters $\mathcal{N}^{t}(0.0,1.0,2.0)$ for the remaining codes. On the other hand, we perturb the StyleGan2 selected groups with parameters: $\mathcal{N}^{t}(0.0,0.2,1.0)$; $\mathcal{N}^{t}(0.0,0.1,1.0)$; $\mathcal{N}^{t}(0.0,0.8,1.0)$; $\mathcal{N}^{t}(0.0,0.8,1.0)$, respectively. Similar to baselines, the learned perturbations are decided by monitoring the walker checkpoints. In general, they follow the same perturbation magnitudes as the random counterpart.
All linear classifiers used for BigBiGan evaluation are trained for $60$ epochs with a batch size of $256$, SGD optimizer, and a learning rate of $30.0$ with cosine decay. StyleGan classifiers follow the same setup, but they are trained for $100$ epochs. For Pascal VOC detection, the R50505050-C4444 backbone is fine-tuned for $24000$ iterations on trainval07+12 split and evaluated on test07.
@@
All our experiments use FFCV (Leclerc et al., 2023) library for efficient data storage and fast loading. ImageNet-$1$K’s images are stored at $256\times 256$ resolution and resized to $128\times 128$ during loading, to match the output resolution of BigBiGan. Regarding LSUN Cars/StyleGan2, instructions to download the $893$K training images can be found at (Karras & Hellsten, 2021). These are $512\times 384$ resolution images, which are stored at $512\times 512$ with black padding, to match StyleGan2 outputs. During loading, images are first center cropped at $384\times 384$, removing padding, and then resized at $128\times 128$. The same preprocessing is applied to the generated images.
@@
Most experiments have been run using $4$ NVIDIA A100100100100-SXM4444-$40$GB GPUs, with an exception for the StyleGan2-trained encoders, which required $8$ GPUs of the same type, due to the larger number of parameters of StyleGan2 w.r.t. BigBiGan. Other minor experiments, like the training of perturbation functions, required only $1$ GPU. To ensure reproducibility, random seeds have always been fixed. For Continuous Sampling, seeds are changed at every iteration and device-specific. This procedure avoids reproducing the same batches during training but allows for consistency throughout different runs.
@@
In BigBiGan experiments, perturbations in the latent space are applied by summing the noise vector to the selected latents, as described in Section 3 of the main paper. For StyleGan2, the final latent space is known as $\mathcal{W}$, and a mapping network is used to perform $f(\mathbf{z})=\mathbf{w}$. Here, $f$ is the mapping network, $\mathbf{w}$ the random latent vector in the $\mathcal{W}$ space, and $\mathbf{z}$ the initial random vector sampled from a truncated Gaussian distribution ($\mathcal{Z}$ space). The positive views in this case are obtained as $f(T_{\mathbf{z}}(\mathbf{z}))=\mathbf{w}$, where $T_{\mathbf{z}}$ is a random or learned perturbation that affects only the selected variables.
@@ Appendix E Ablation Studies
Previous studies (Jahanian et al., 2021; Li et al., 2022) show that the random and learned baselines still benefit from the application of $T_{\mathbf{x}}$ SimCLR augmentations on top of the generated views. In detail, they perform some ablation studies training encoders on Imagenet-$100$ and a subset of LSUN cars, where different combinations of pixel-space augmentations are tested: cropping and horizontal flipping, grayscale and color jittering, none of the previous or all of them. When utilizing ML perturbations, we qualitatively observe that the generated views can assume various realistic colors (as seen for example in Figure 4 of the main paper). For this reason, in the experimental analysis, we only use cropping and flipping as pixel space augmentations.
To validate our hypothesis, in Figures 7 and 8 we report the result of the same experiments of previous work, but performed on our ML views. In all cases, removing color jittering and grayscale operations increases the final linear evaluation Top-1 accuracy. These observations empirically confirm that ML views do not need additional color transformations.
@@ Appendix F NVAE Experiments
In the main paper, experiments are conducted on two state-of-the-art GAN methods. To showcase the general properties of the proposed approach, we conducted an additional ablation study where we trained a small NVAE (Vahdat & Kautz, 2020) with $3$ latent scales on Cifar10101010 (Krizhevsky et al., 2009) ($32\times 32$ resolution). To test its abilities as a data source, we replicated the experiments of Table 2 for the random case, employing the SimSiam framework and Continuous Sampling. Table 4 shows the final Top-$1$ and Top-$5$ accuracies obtained on the Cifar10101010 evaluation set, where the parameters for sampling anchors have been learned during training of the NVAE. (a): three different perturbation magnitudes for the baseline have been considered. (b) The first two latents are fixed, and latent $2$ is sampled twice (anchor and positive). The fine-grained control offered by the multiple variables achieves better scores w.r.t. the single latent baseline also when NVAE is used, demonstrating the general properties of the proposed approach.
Run $T_{\mathbf{z}}$ Top-1 Acc Top-5 Acc
random $N^{t}(0.,0.05,2.)$ $40.20$ $88.05$
(a) random $N^{t}(0.,0.10,2.)$ $42.09$ $88.88$
random $N^{t}(0.,0.15,2.)$ $40.15$ $88.19$
(b) ML random code $2$ resample $\mathbf{43.86}$ $\mathbf{89.69}$
Table 4: Evaluation on Cifar10101010 (Krizhevsky et al., 2009) of encoders trained using an NVAE (Vahdat & Kautz, 2020) with $3$ latent codes as a data source. (a) Positive views are obtained with random perturbation, as in (Jahanian et al., 2021). (b) positive views are obtained by resampling and substitution of the last code only, leaving the first two variables unchanged.
@@ Appendix G Transfer Learning
Table 5 contains the results for each transfer classification learning experiment performed on top of the SimSiam pre-trained encoders using BigBiGan / ImageNet-$1$K. The results refer to the last column of Table 2, where only the mean Top-$1$ accuracy over the $7$ target datasets has been reported. For a better comparison of encoders’ generalization capabilities, each run was tested with $5$ different seeds, and the mean Top-$1$ accuracy was taken.
On each dataset, the results are computed using the test set where available, otherwise on the validation set, maintaining the original splits. For DTD (Cimpoi et al., 2014) the first split between the proposed ones has been employed, while for Caltech101101101101 (Fei-Fei et al., 2004) we selected a random split of $30$ train images per class, using the remaining for testing. In this case, all the background images (distractors) have been removed.
Encoder Top-$1$ Accuracy on Target Dataset
Birdsnap Caltech101101101101 Cifar100100100100 DTD Flowers102102102102 Food101101101101 Pets
@@
Table 5: Transfer classification Top-$1$ accuracy’s results for each pre-trained encoder using the SimSiam framework and BigBiGan/ImageNet-$1$K as a data source. $7$ different target datasets have been tested.
@@ Appendix H Qualitative Visualizations
Figure 9 and Figure 10 show some examples of generated views using BigBiGan and StyleGan2, respectively. In each Figure, rows represent different views, and columns display (from left to right): the initial anchor image, random vs. learned baselines, and the views obtained with our method. All the examples are generated using the same hyperparameters, which were also used for training the contrastive encoders. In order to allow a better comparison of the different methods, no further $T_{\mathbf{x}}$ augmentations are applied on top of the generated images. As depicted in both figures, our generated views can produce a wide range of transformations, including realistic color changes that allow us to eliminate the color jittering $T_{\mathbf{x}}$ augmentation.
Figure 9: Example of generated views from BigBiGan anchors. From left to right: anchor, random and learned baseline views, random and learned HL views.
Figure 10: Example of generated views from StyleGan2 anchors. From left to right: anchor, random and learned baseline views, random and learned HL views.

### v2

@@ 4.3 Continuous Sampling
All our encoders are trained using Continuous Sampling, except for SimCLR, which follows the previous setup. Additionally, to compare overall training speed to standard data loading, we trained a ResNet-$18$ model with SimCLR for $20$ epochs on ImageNet-$100$, on $4$ NVIDIA A100-SXM4-40GB GPUs and different batch sizes ($32\times 4$, $64\times 4$, $128\times 4$, $256\times 4$). The experiment compares the standard PyTorch Paszke et al. (2019) loader, the efficient FFCV Leclerc et al. (2023) loader (both with $8$ workers), and the BigBiGan generator. Figure˜5 displays our findings, reporting the mean GPU seconds per epoch, the $\text{CO}_{2}$ emissions rate, and the total $\text{CO}_{2}$ emissions estimated using CodeCarbon Schmidt et al. (2021). Continuous Sampling proves significantly faster than torch loader and only marginally slower than FFCV. In terms of $\text{CO}_{2}$ emissions rate, the use of BigBiGan led to higher energy consumption, due to intensive GPU usage. Nevertheless, in terms of total $\text{CO}_{2}$ the values remain comparable with torch. In conclusion, the increasing efficiency and precision of modern image generation models, especially fast-sampling GANs, make Continuous Sampling an interesting alternative to conventional data-loading techniques, allowing great image variability while maintaining comparable training times.
@@ 5 Discussion and Conclusions
Our work showcases MLVGMs as a distinct category of models, offering new tools to assess the impact of latent variables. Specifically, the proposed Monte Carlo quantification method supports previous empirical observations on the “global-to-local” nature of MLVGMs, but allows a more in-depth and quantitative analysis. Nonetheless, our algorithm remains limited by the nature of Mutual Information itself, which is invariant under invertible transformations. As a consequence, some small perturbations (e.g. a constant scaling to each pixel), will be approximated as a 0 shift by our method. However, since the differences in average perturbation between each latent space are usually significant, it is reasonable to assume that such small approximations do not invalidate our findings. More specifically, we reveal that modern gan-based MLVGMs, such as BigGan and StyleGan employ over or under-utilized variables in the generative process, setting up a base for possible architectural improvements. In terms of view generation, our method has proven its superiority, surpassing previous perturbation strategies applied to single-variable models. However, it does not address the inherent challenge of SSCRL: views are defined upon “reasonable” thresholds, since “optimal” positives depend on the specific downstream task. Regarding generative models as a data source, they offer potential solutions to issues associated with real datasets, such as privacy concerns and usage rights Kaissis et al. (2020); DuMont Schütte et al. (2021). However, generative models can inherit biases from the original data Asim et al. (2020), and latent perturbations may amplify them, propagating to downstream models. Therefore, techniques to mitigate these biases could be considered Tan et al. (2020); Teo et al. (2023). As for Continuous Sampling, while it reduces disk usage and data diversity, scaling to high image resolutions or large models may be very GPU-intensive and result in high CO2 emissions, limiting usability.
@@ Appendix B Training Dynamics
As explained in Section˜3.1, we train each perturbation function $T_{\mathbf{z}}^{i}$ separately, and initialize as the identity function. This allows InfoNCE loss in Equation˜5 to start from a very low value. During training, $T_{\mathbf{z}}^{i}$ progressively enhances diversity in the views, reducing $I(\mathbf{X};\mathbf{X}^{\prime})$, and causing InfoNCE to increase over time. An example of these training dynamics is illustrated in Figure˜6(a). Figure˜6(b) shows some qualitative examples of how views in the pixel space evolve during training. Specifically, the image on the left represents the anchor, while other images show the corresponding perturbed view as the magnitude increases.
@@ Appendix C Continuous Sampling
We detail the batch generation procedure for anchors and positive views in Algorithm˜1. As explained in the main paper, the algorithm requires loading the pre-trained generative model (MLVGM in our case) on the same GPU device as the encoder model to be trained. In this way, the generated batch will be already on the correct device, eliminating the need for data loading.
@@
[…]
@@
4:  $A_{b}\leftarrow\emptyset$ $\triangleright$ initialize empty set of anchor variables
5:  $P_{b}\leftarrow\emptyset$ $\triangleright$ initialize empty set of positive variables
@@
7:   $a_{i}\sim N^{t}(\mu,\sigma,t)$ $\triangleright$ sample anchor code
8:   $p_{i}\leftarrow T_{\mathbf{z}}^{i}(a_{i})$ $\triangleright$ generate positive code
@@
12:  $\mathbf{a}_{b}\leftarrow g(A_{b})$ $\triangleright$ generate anchor image
13:  $\mathbf{p}_{b}\leftarrow g(P_{b})$ $\triangleright$ generate positive image
@@
17:return $\mathcal{A},\mathcal{P}$ $\triangleright$ SSCRL Encoder’s inputs
@@ Appendix D View Generation Training Details
For sampling positives, the random baselines are trained with $\mathcal{N}^{t}(0.0,0.2,2.0)$ and $\mathcal{N}^{t}(0.0,0.25,0.9)$ for each generator, respectively BigBiGan and StyleGan2. Following the suggestions of the original work (Li et al., 2022), we monitor the generated views and select the best checkpoints for the learned baselines. For our ML views, BigBiGan samples positives by fixing the first latent and with parameters $\mathcal{N}^{t}(0.0,1.0,2.0)$ for the remaining codes (including for simplicity the unused latent). On the other hand, we perturb the StyleGan2 selected groups with parameters: $\mathcal{N}^{t}(0.0,0.2,1.0)$; $\mathcal{N}^{t}(0.0,0.1,1.0)$; $\mathcal{N}^{t}(0.0,0.8,1.0)$; $\mathcal{N}^{t}(0.0,0.8,1.0)$, respectively. Similar to baselines, the learned perturbations are decided by monitoring checkpoints. In general, they follow the same perturbation magnitudes as the random counterpart.
All linear classifiers used for BigBiGan evaluation are trained for $60$ epochs with a batch size of $256$, SGD optimizer, and a learning rate of $30.0$ with cosine decay. StyleGan classifiers follow the same setup, but they are trained for $100$ epochs. For Pascal VOC detection, the R5050-C44 backbone is fine-tuned for $24000$ iterations on trainval07+12 split and evaluated on test07.
@@
All our experiments use FFCV (Leclerc et al., 2023) library for efficient data storage and fast loading. ImageNet-11K’s images are stored at $256\times 256$ resolution and resized to $128\times 128$ during loading, to match the output resolution of BigBiGan. Regarding LSUN Cars/StyleGan2, instructions to download the 893893K training images can be found at (Karras & Hellsten, 2021). These are $512\times 384$ resolution images, which are stored at $512\times 512$ with black padding, to match StyleGan2 outputs. During loading, images are first center cropped at $384\times 384$, removing padding, and then resized at $128\times 128$. The same preprocessing is applied to the generated images.
@@
Most experiments have been run using $4$ NVIDIA A100100-SXM44-4040GB GPUs, with an exception for the StyleGan2-trained encoders, which required $8$ GPUs of the same type, due to the larger number of parameters of StyleGan2 w.r.t. BigBiGan. Other minor experiments, like the training of perturbation functions, required only $1$ GPU. To ensure reproducibility, random seeds have always been fixed. For Continuous Sampling, seeds are changed at every iteration and device-specific. This procedure avoids reproducing the same batches during training but allows for consistency throughout different runs.
@@
In BigBiGan experiments, perturbations in the latent space are applied by summing the noise vector to the selected latents, as described in Section˜3 of the main paper. For StyleGan2, the final latent space is known as $\mathcal{W}$, and a mapping network is used to perform $f(\mathbf{z})=\mathbf{w}$. Here, $f$ is the mapping network, $\mathbf{w}$ the random latent vector in the $\mathcal{W}$ space, and $\mathbf{z}$ the initial random vector sampled from a truncated Gaussian distribution ($\mathcal{Z}$ space). The positive views in this case are obtained as $f(T_{\mathbf{z}}(\mathbf{z}))=\mathbf{w}$, where $T_{\mathbf{z}}$ is a random or learned perturbation that affects only the selected variables.
@@ Appendix E Ablation Studies
Previous studies (Jahanian et al., 2021; Li et al., 2022) show that the random and learned baselines still benefit from the application of $T_{\mathbf{x}}$ SimCLR augmentations on top of the generated views. In detail, they perform some ablation studies training encoders on Imagenet-$100$ and a subset of LSUN cars, where different combinations of pixel-space augmentations are tested: cropping and horizontal flipping, grayscale and color jittering, none of the previous or all of them. When utilizing ML perturbations, we qualitatively observe that the generated views can assume various realistic colors (as seen for example in Figure˜4 of the main paper). For this reason, in the experimental analysis, we only use cropping and flipping as pixel space augmentations.
To validate our hypothesis, in Figures˜7 and 8 we report the result of the same experiments of previous work, but performed on our ML views. In all cases, removing color jittering and grayscale operations increases the final linear evaluation Top-1 accuracy. These observations empirically confirm that ML views do not need additional color transformations.
@@ Appendix F NVAE Experiments
In the main paper, experiments are conducted on two state-of-the-art GAN methods. To showcase the general properties of the proposed approach, we conducted an additional ablation study where we trained a small NVAE (Vahdat & Kautz, 2020) with $24$ latent variables, organized in $3$ groups, on Cifar1010 (Krizhevsky et al., 2009) ($32\times 32$ resolution). The model achieves a FID of $23.63$, and an IS score of $6.36\pm 0.16$.
At first, we measure the impact of each latent group replicating the experiments of Table˜1. The results, included in Table˜4, confirm that the “global-to-local” trend is valid also in this case. Moreover, we observe that the first group is highly influencing the final outcome with respect to the remaining groups.
Table 4: Results of the MC simulation on the NVAE model pre-trained on Cifar-10. For each latent group $g$ we show the final InfoNCE loss value and the estimated mean ($\mu$) and standard deviation ($\sigma$) of the corresponding inferred distribution.
latent group loss estimated $q^{i}$
($g$) (InfoNCE) ($\mu$) ($\sigma$)
11 $0.60$ $0.04$ 1e31e-3
22 $0.60$ $0.96$ $0.41$
33 $0.60$ $1.18$ $1.25$
Afterwards, we test the model’s abilities as a data source, replicating the experiments of Table˜2 for the random case with SimSiam framework and Continuous Sampling. Table˜5 shows the final Top-$1$ and Top-$5$ accuracies obtained on the Cifar1010 evaluation set. In this case, the absolute best result is achieved by the same encoder trained on real data ($79.15\%$ accuracy), with a large gap ($54.95\%$) when using NVAE as a data source without further latent perturbations. Adding the same perturbations to all groups provides an immediate accuracy boost ($5\%$), with various standard deviations. After observing the results in Table˜4, we decided to fix the first group (overused), and find that applying larger perturbations to groups 2-3 only allows for a larger std and additional boost, with a final accuracy of $60.75\%$. These experiments prove the feasibility of our methods also on Autoencoder-based MLVGMs, despite the small model employed.
Table 5: Evaluation on Cifar1010 (Krizhevsky et al., 2009) of encoders trained using an NVAE (Vahdat & Kautz, 2020) with $3$ latent groups as a data source.
Encoder $T_{\mathbf{z}}$ Top-1 Accuracy (%)
Baseline real - $79.15$
Baseline synth - $54.95$
Random $N^{t}(0.,0.2,1.)$ $59.44$
Random $N^{t}(0.,0.3,1.)$ $59.61$
Random $N^{t}(0.,0.4,1.)$ $59.34$
ML Random $N^{t}(0.,0.3,1.)$ $59.21$
ML Random $N^{t}(0.,0.4,1.)$ $60.75$
ML Random $N^{t}(0.,0.5,1.)$ $59.98$
@@ Appendix G Transfer Learning
Table˜6 contains the results for each transfer classification learning experiment performed on top of the SimSiam pre-trained encoders using BigBiGan / ImageNet-11K. The results refer to the last column of Table˜2, where only the mean Top-$1$ accuracy over the $7$ target datasets has been reported. For a better comparison of encoders’ generalization capabilities, each run was tested with $5$ different seeds, and the mean Top-$1$ accuracy was taken.
On each dataset, the results are computed using the test set where available, otherwise on the validation set, maintaining the original splits. For DTD (Cimpoi et al., 2014) the first split between the proposed ones has been employed, while for Caltech101101 (Fei-Fei et al., 2004) we selected a random split of $30$ train images per class, using the remaining for testing. In this case, all the background images (distractors) have been removed.
Encoder Top-11 Accuracy on Target Dataset
Birdsnap Caltech101101 Cifar100100 DTD Flowers102102 Food101101 Pets
@@
Table 6: Transfer classification Top-11 accuracy’s results for each pre-trained encoder using the SimSiam framework and BigBiGan/ImageNet-11K as a data source. 77 different target datasets have been tested.
@@ Appendix H Qualitative Visualizations
Figure˜9 and Figure˜10 show some examples of generated views using BigBiGan and StyleGan2, respectively. In each Figure, rows represent different views, and columns display (from left to right): the initial anchor image, random vs. learned baselines, and the views obtained with our method. All the examples are generated using the same hyperparameters, which were also used for training the contrastive encoders. In order to allow a better comparison of the different methods, no further $T_{\mathbf{x}}$ augmentations are applied on top of the generated images. As depicted in both figures, our generated views can produce a wide range of transformations, including realistic color changes that allow us to eliminate the color jittering $T_{\mathbf{x}}$ augmentation.
Figure 9: Example of generated views from BigBiGan anchors. From left to right: anchor, random and learned baseline views, random and learned ML views.
Figure 10: Example of generated views from StyleGan2 anchors. From left to right: anchor, random and learned baseline views, random and learned ML views.